OpenStudy (thatcreepyboy):

A weather balloon holds 2,850 cubic meters of helium. The density of helium is 0.1685 kilograms per cubic meter. How many kilograms of helium does the balloon contain? Please round your answer to the nearest whole number. Only enter a numerical value in the answer blank.

OpenStudy (mrnood):

the question tells you that EACH cu metre weighs 0.1685kg so multiply that by the number of cu meteres in the balloon for your answer

OpenStudy (imstuck):

set up a ratio/proportion/equivalency situation where what you are solving for is kilograms. I teach it like this so it's easy to see how things cancel each other out, etc. \[2850m ^{3}\times \frac{ .1685k }{ 1m ^{3} }\]

OpenStudy (imstuck):

By setting it up like that, you have an equivalency going on between kilograms and cubic meters. The problem tells you that 1 cubic meter is equal to .1685 kg. Now the cubic meters cancel each other out, leaving you with 2850 times .1685 with your only label remaining being kg, which is what you're looking for.

OpenStudy (thatcreepyboy):

Do I multiply 2,850 by 1685?

OpenStudy (thatcreepyboy):

@MrNood @IMStuck

OpenStudy (thatcreepyboy):

Idk....

OpenStudy (welshfella):

density = mass / volume so plugging in the given values 0.1685 = M / 2850 now multiply left side by 2850 will give you the mass M.

OpenStudy (welshfella):

M = 0.1685 * 2850 use your calculator

OpenStudy (thatcreepyboy):

I got 480.225.

OpenStudy (thatcreepyboy):

Is it right? Or am I doing something wrong?.

ganeshie8 (ganeshie8):

Yes, thats right. You should trust yourself more :)

ganeshie8 (ganeshie8):

Hey they want you round the answer to the nearest whole number

ganeshie8 (ganeshie8):

What is 480.225 rounded to the nearest whole number ?

OpenStudy (thatcreepyboy):

480?

ganeshie8 (ganeshie8):

Looks good!

OpenStudy (thatcreepyboy):

Thank you! :D